OpenStudy (anonymous):

Solve the system -3x+y+3z=-2 x+7y-2z=5 4x+6y-5z=7 Please help

OpenStudy (anonymous):

in what way you want to solve it?

OpenStudy (anonymous):

If it is consistent, inconsistent, or dependent.

OpenStudy (anonymous):

Okay, I have found it to be dependent but how do I find x and y?

OpenStudy (anonymous):

let z=k and write first two equations in terms of x and y and k then find x and y in terms of k.

OpenStudy (anonymous):

What is k?

OpenStudy (anonymous):

k is an arbitrary constant.

OpenStudy (anonymous):

So if I was to use x+7y-2z=5 to find x how would I do it?

OpenStudy (anonymous):

then you will find x and y in terms of z.

OpenStudy (anonymous):

I still don't get it

hero (hero):

Split the system into pairs. Then eliminate the same variable from each pair. Afterwards what you will end up with is a system of two equations, which you should be able to solve.

OpenStudy (anonymous):

I did that and 0=0 but I still have to find x and y

hero (hero):

a) -3x+y+3z=-2 b) x+7y-2z=5 c) x+7y-2z=5 d) 4x+6y-5z=7

OpenStudy (anonymous):

Right and I ended up with -22y+3z=-13 22y-3z=13

hero (hero):

Multiply both sides of equation b) by -3 Multiply both sides of equation c) by 4 a) -3x+y+3z=-2 b) -3x-21y +6z=-15 c) 4x+28y-8z=20 d) 4x+6y-5z=7 Then eliminate the x variable to get: 22y -3z = 13 22y -3z = 13

OpenStudy (anonymous):

Was I wrong to multiply by 3 and -4?

hero (hero):

Not necessarily, but you should have ended up with what I got.

OpenStudy (anonymous):

I did. But I ended up with the equations canceling out and 0=0

hero (hero):

Nevermind. It really doesn't matter. Either way, you end up with 0 = 0

hero (hero):

Basically, it is not possible to solve for y or z.

OpenStudy (anonymous):

Well, my assignment says if the answer is dependent I have to express the solution in terms of one of the variables.

OpenStudy (anonymous):

It is supposed to look like this {(numbers, numbers, variable)|(epsilon)R}

hero (hero):

Try using matrices to figure it out

OpenStudy (anonymous):

What?

hero (hero):

If you've never heard of matrices, then disregard

OpenStudy (anonymous):

-3x+y=-2-3z ...(1) x+7y=5+2k ...(2) multiply (2) by 3 and add in (1) 22y=13+3ky \[y=\frac{ 13+3k }{ 22 }\] similarly x=? solution is x=? \[y=\frac{ 13+3k }{ 22 }\] z=k giving different values to k we get different values of x,y,z.